OpenStudy (anonymous):

PLZZZZZZZZZZZZZZZZZZZZ HELP ME WTH THIS QUESTION ILL GIVE MEDAL!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!1

OpenStudy (anonymous):

?

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

Solve for c. c plus two-fifths equals negative one-tenth c equals negative one-fifth c equals negative one-half c equals one-tenth c equals three-fifths

OpenStudy (anonymous):

c= -1/2

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

WAIT!

OpenStudy (anonymous):

Solve for g. one-twelfth equals g minus five-sixths g equals negative eleven-twelfths g equals negative three-fourths g equals three-fourths g equals eleven-twelfths

OpenStudy (anonymous):

g=11/12

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

Solve for k. k minus five-twelfths equals negative four-twelfths k equals negative three-fourths k equals negative one-twelfth k equals three-fourths k equals one-twelfth

OpenStudy (anonymous):

k=1/12